Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, January 19, 1999

The following is a brief summary of the environmental monitoring activities that were completed between January 19, 1999 1:00 and 1:00 on January 20, 1999. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

01/19/99 (Tuesday)

Segment 1 (Emidio Pump Station to Copco Road): Report for 19 January 1999: The final weld for Pacific Pipeline was successfully completed this afternoon at the tie in to Emidio Pump Station. Crews began some backfill of the pipe surrounding the weld, that remains to be coated. Other backfill included the nearly complete area at the west I-5/Tejon bore site, as well as the tie in location on Copco Road. The other tie in locations, Flying J, north and south Grapevine Pump Station, were completed today. General clean up occurred throughout the segment. Revegetation efforts continued along Rock Road.

Segment 5 (Lynwood to Long Beach): Report for 1-19-99. After cutting out the valve on the west end of the Edison bore crews excavated several feet of earth and placed gravel on the bottom of the bore pit to level the site. Efforts were then made to place the large concrete valve box into the pit and re-weld the valve into place. Huntway crews welded tube turns and saw caught the roadway between the receiving bore pit site and Lomita St. Further, crews marked out the right of way and organized gear and materials for trenching. Meanwhile, crews within Tosco at the Wilmington station reshot x-rays on several pipes, repaired damaged coating on the sump drain lines and connected the communication line to the systems control center. In addition, excavation of the large bore pit progressed while commission crews examined the station. Segment 4B (Lynwood to El Segundo): Crews continued installing headers and brackets for the roof section at Lynwood station while crews working at El Segundo began painting.

Inspection Summar

In the past 24 hours, 2 inspection event(s) were reported for 2 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page